Marie and Andy were married 21
years ago. Andy was three times
as old as Marie. Today, he is twice
as old as she is. How old was
Marie on her wedding day?

Answer by macston(5194)

.
M=Marie's age 21 years ago; A=Andy's age 21 years ago=3M
.
A+21=2(M+21)
3M+21=2M+42
M=21
ANSWER: Marie was 21 years old on her wedding day.
.
CHECK:
A=3M=3(21)=63
A+21=2(M+21)
63+21=2(21+21)
84=2(42)
84=84